# From Slices to Structures: Unsupervised 3D Reconstruction of Female Pelvic Anatomy from Freehand Transvaginal Ultrasound

### v1

@@ Abstract
Volumetric ultrasound has the potential to significantly improve diagnostic accuracy and clinical decision-making, yet its widespread adoption remains limited by dependence on specialized hardware and restrictive acquisition protocols. In this work, we present a novel unsupervised framework for reconstructing 3D anatomical structures from freehand 2D transvaginal ultrasound (TVS) sweeps, without requiring external tracking or learned pose estimators. Our method adapts the principles of Gaussian Splatting to the domain of ultrasound, introducing a slice-aware, differentiable rasterizer tailored to the unique physics and geometry of ultrasound imaging. We model anatomy as a collection of anisotropic 3D Gaussians and optimize their parameters directly from image-level supervision, leveraging sensorless probe motion estimation and domain-specific geometric priors. The result is a compact, flexible, and memory-efficient volumetric representation that captures anatomical detail with high spatial fidelity. This work demonstrates that accurate 3D reconstruction from 2D ultrasound images can be achieved through purely computational means, offering a scalable alternative to conventional 3D systems and enabling new opportunities for AI-assisted analysis and diagnosis.
{IEEEkeywords}
Unsupervised 2D-to-3D reconstruction, transvaginal ultrasound, Gaussian Splatting, ultrasound rasterization
@@ 1 Introduction
Three-dimensional (3D) ultrasound (US) imaging plays an essential role in obstetrics and gynecology, providing clinicians with a spatial context that is crucial for diagnosis, surgical planning, and treatment monitoring [1, 2]. Transvaginal ultrasound (TVS) is widely used for high-resolution imaging of pelvic organs due to its affordability, safety, and minimally invasive nature, extending 2D TVS to 3D imaging remains technically challenging. Existing 3D US systems typically rely on external tracking mechanisms or mechanically swept probes. However, while 3D ultrasound imaging is commonly performed both transabdominally and transvaginally, especially in obstetric applications such as fetal imaging, most commercially available 3D systems are primarily designed for abdominal or general freehand use and are not specifically optimized for transvaginal applications in gynecological contexts [3, 4]. In addition to being physically cumbersome and difficult to integrate into routine clinical workflows, these systems often suffer from lower resolution compared to conventional 2D imaging. The reliance on mechanical scanning or external tracking can introduce motion artifacts, particularly in dynamic or less controlled scanning environments. Furthermore, the need for specialized hardware increases system complexity and cost, which hinder widespread adoption.
@@
In this work, we propose an alternative paradigm: an unsupervised method for 3D US reconstruction that is free from the need for any external tracking or learned pose estimation. Our method is specifically designed for the gynecological domain and operates on standard 2D TVS sweeps acquired freehand. It leverages recent advances in volumetric rendering, adapting Gaussian Splatting [9] to the constraints and physics of US imaging. Our approach uses sensorless geometric estimation of probe motion to establish spatial coherence between slices, and reconstructs anatomical structures using anisotropic Gaussian primitives optimized directly from US data.
Our method consists of three key components: (1) sensorless probe position estimation, based on sweep consistency and domain-specific geometric assumptions, (2) a volumetric representation using learned isotropic Gaussians that model anatomy with flexibility and memory efficiency, and (3) a custom CUDA-optimized differentiable rasterizer tailored to the non-optical, slice-based nature of US acquisition. Unlike voxel-based interpolation or view synthesis, Gaussian Splatting enables detailed and adaptive reconstruction without the constraints of a fixed voxel grid. Importantly, our method makes no assumptions about trajectories, requires no calibration data, and does not rely on training with large datasets.
@@
  • A CUDA-optimized differentiable rasterizer tailored for discrete and anisotropic slice-based rendering.
@@ 2 Related Work
Early freehand US systems often relied on external tracking hardware, such as electromagnetic sensors, to determine the spatial position and orientation of each 2D slice [10]. However, these systems are not suitable for TVS due to their bulk, line-of-sight limitations, and added patient discomfort [5]. To overcome these constraints, sensorless reconstruction methods were developed based on signal processing techniques. One foundational approach used speckle decorrelation, where the characteristic speckle pattern in US images was leveraged to infer probe movement [11, 12]. Despite its innovation, this method is limited by noise sensitivity and an inability to handle complex multi-axis motion, especially common in freehand TVS sweeps [13]. Spatial correlation techniques, such as normalized cross-correlation or sum-of-absolute-differences, were also explored to estimate small in-plane shifts [14]. These methods, while effective for minor linear movements, performed poorly with rotational or elevational displacement. More sophisticated approaches attempted to improve robustness through statistical models and trajectory smoothing, such as graph-based optimization constrained by speckle variation [15]. Nevertheless, these approaches remained fragile and were rarely adopted outside of research settings.
@@
With the rise of deep learning, supervised learning methods for pose estimation gained popularity. These approaches train convolutional or recurrent neural networks to predict either inter-frame transformations or directly reconstruct 3D volumes. Early models such as [6] demonstrated that CNNs could infer relative probe motion from US frames, enabling trackerless 3D reconstruction. Later methods like $\text{DC}^{2}$-Net [16] introduced contrastive and temporal losses to improve trajectory consistency across sequences. Another noteworthy model is PLPPI, which incorporates domain-specific physics constraints into the learning pipeline, enabling improved handling of out-of-plane motion [17]. However, a shared limitation of these models is their dependency on ground truth data from tracked acquisitions for training and evaluation. In many clinical workflows—particularly for TVS—such tracking data is unavailable or infeasible to collect. Some recent works like RecON [18] explore online learning, updating models in real-time as new frames are acquired. While promising for reducing drift and latency, these methods still presuppose access to prior 3D volumes or strong priors, making them unsuitable for fully unsupervised clinical settings.
@@
In parallel, implicit representations like Neural Radiance Fields (NeRF) [19] and subsequent developments (e.g., ImplicitVol [4]) model 3D structures as continuous functions learned by neural networks. These models offer resolution independence and are memory-efficient but require extensive training and typically rely on known camera poses—conditions not met in US. These models offer resolution independence and are memory-efficient but typically require extensive training and known camera poses—conditions that are generally not met in US unless a tracking device is present. Hybrid methods such as RapidVol [7] attempt to blend implicit and explicit approaches, using tri-planar maps refined by lightweight networks to accelerate training and inference. However, even these models typically assume access to 3D volumes for training or initialization.
@@
Beyond motion estimation, the choice of volumetric representation plays a central role in reconstruction quality and efficiency. Traditional explicit methods such as voxel grids offer high interpretability and easy integration into existing medical imaging pipelines. However, they suffer from cubic memory growth with resolution and are thus computationally intensive for fine anatomical detail. More recently, techniques like Gaussian Splatting have emerged as a powerful alternative. Originally developed for real-time rendering in computer vision [9], Gaussian Splatting represents scenes using parameterized Gaussian primitives whose locations, shapes, and radiance are optimized to match rendered views. UltraGauss [8] adapted this to fetal US imaging by incorporating anisotropic Gaussians and integrating pose predictions from QAERTS [20]. This approach improves rendering quality but still depends on learned pose models, which hinders generalization and deployment in tracking-free environments.
@@
In contrast to these existing approaches, our method introduces the first application of Gaussian Splatting for 3D US that is entirely unsupervised and sensorless. We eliminate the need for pose estimation models, supervision, or tracking hardware by leveraging geometric assumptions on sweep motion and optimizing isotropic Gaussian primitives directly from 2D TVS slices. This enables anatomically accurate, scalable 3D reconstruction using only standard 2D US equipment, significantly lowering the barrier to clinical deployment and advancing the capabilities of low-cost US imaging.
@@ 3.1 Synthetic Data
Figure 2: Synthetic ultrasound slices generated from a simulated uterus model using a renderer. Left: Sagittal view showing longitudinal anatomy. Right: Transversal view perpendicular to the sagittal plane.
@@ 3.2 Real Data
12 TVS sweeps were acquired from 6 patients (2 sweeps per patient) using a GEHC Voluson Expert 22 ultrasound machine. Table 1 summarizes the main parameters of the data. Each sweep represents either a sagittal or transversal view, with probe motion approximated as a smooth angular trajectory. An example image for both modalities can be seen in Fig. 3. Each patient provided written informed consent prior to participation, and all procedures adhered to strict ethical standards to ensure confidentiality and privacy. Since real data lacks tracked probe poses, we assume consistent frame intervals and smooth probe motion. We initialize the pose of each frame using equal angular spacing. This assumption mirrors the angular sampling strategy used in generating synthetic data. Other than resizing and padding images when needed, no additional transformations or data augmentation techniques are applied, ensuring that the intrinsic properties of the images are preserved. Our method assumes (i) temporally ordered slices with approximately uniform angular spacing, and (ii) minimal tissue deformation during the sweep. These assumptions are critical for effective sensorless reconstruction, and we analyze their impact in Section 4.
@@
fps 10-20
Duration 10 seconds
@@
Age 38.33 (8.48)
Height [m] 1.65 (0.04)
Weight [kg] 64.53 (5.32)
@@ 4.1 Overview: From Gaussian Splatting to Ultrasound Imaging
However, US imaging presents fundamental differences. Each image (or “slice”) corresponds to a narrow, planar cross-section of tissue, with limited or no overlap between neighboring slices. As a result, there is no redundancy between views to help maintain scene consistency. Additionally, the physics of US introduces directionality and discontinuities that standard perspective projection does not capture. This stands in contrast to Gaussian Splatting for optical images, where we work with projections of the scene onto a camera sensor, inherently providing depth information. Since US does not have this projection component and instead captures direct cross-sections, approaches relying on projection geometry cannot be directly applied. Misaligned Gaussian primitives—especially in depth—can significantly affect the rendering, since each slice is extremely thin. To address these challenges, we adapt Gaussian Splatting, building a slice-aware differentiable rasterizer specifically designed for US data. Our framework accounts for the sparse and anisotropic nature of slices while retaining the strengths of the Gaussian formulation—compact, continuous scene representation and differentiability.
@@ 4.2 Slice-Based Differentiable Rasterization
Our rasterizer computes the intensity of each pixel by aggregating contributions from a set of 3D Gaussian primitives. We begin by describing the formulation in terms of its constituent steps. Each Gaussian $g$ is defined by a set of parameters:
@@
  • $\Sigma_{g}\in\mathbb{R}^{3\times 3}$: the covariance matrix describing its shape and orientation. It is factored as $\Sigma_{g}=(S_{g}R_{g})^{\top}(S_{g}R_{g})$, where $S_{g}$ is a diagonal scale matrix and $R_{g}$ is a rotation matrix derived from a unit quaternion,
@@
Each pixel $p$ in an US slice is associated with a known 3D coordinate $\mathbf{c}_{p}\in\mathbb{R}^{3}$, obtained via slice pose and pixel location. For each Gaussian $g$, we compute the offset vector from the Gaussian center to the pixel coordinate:
@@
This formulation defines each pixel as the sum of soft and spatially-varying influences from 3D Gaussians. By using Gaussian functions instead of hard surface projections, the method can represent continuous, soft-tissue-like structures and remains differentiable. It is especially well-suited for sparse and anisotropic data such as US slices, where direct projection is unreliable and information across slices is limited.
4.3 Gradient-Based Optimization
To enable end-to-end training of our differentiable rasterizer, we must compute gradients of the loss $L$ with respect to all learnable parameters of each Gaussian primitive. These include the intensity $I_{g}$, opacity $o_{g}$, mean position $\mathbf{\mu}_{g}$, and covariance $\Sigma_{g}$ (parameterized via scale and rotation). Unlike standard neural network layers, our rasterizer is implemented as a custom CUDA kernel for performance reasons, which precludes the use of automatic differentiation frameworks such as PyTorch’s autograd. Therefore, we manually derive analytic gradients for all parameters using the chain rule and matrix calculus. The key components of this gradient-based optimization pipeline are as follows:
  • Intensity and opacity gradients $\left(\frac{\partial L}{\partial I_{g}},\frac{\partial L}{\partial o_{g}}\right)$: These are computed directly from the contribution term $\alpha_{gp}=o_{g}\exp(e_{gp})$, where each Gaussian’s influence on a pixel is determined by how close it is and how opaque it is.
  • Mean position gradients $\left(\frac{\partial L}{\partial\mathbf{\mu}_{g}}\right)$: The spatial location of the Gaussian strongly affects the rendering. We compute how shifting the Gaussian in 3D space changes its contribution to all nearby pixels.
  • Covariance gradients $\left(\frac{\partial L}{\partial\Sigma_{g}}\right)$: The shape and spread of the Gaussian control its spatial footprint. We derive the gradient of the Mahalanobis exponent and backpropagate through the decomposition $\Sigma_{g}=(S_{g}R_{g})^{\top}(S_{g}R_{g})$.
  • Scale and rotation gradients: Since we do not directly optimize $\Sigma_{g}$, we propagate the covariance gradient to scale and rotation components. The scale is assumed diagonal, while rotation is parameterized via unit quaternions.
  • Quaternion backpropagation: We compute gradients through the rotation matrix $R(q_{g})$ and its analytical Jacobian with respect to the quaternion $q_{g}$. We also account for the normalization of quaternions to ensure they remain on the unit sphere.
This modular decomposition of the backward pass ensures numerical stability, supports high-performance GPU execution, and allows all Gaussian parameters to be learned efficiently during training. For full derivations, see Sec. .1.

### v2

@@ Abstract
Volumetric ultrasound has the potential to significantly improve diagnostic accuracy and clinical decision-making, yet its widespread adoption remains limited by dependence on specialized hardware and restrictive acquisition protocols. In this work, we present a novel unsupervised framework for reconstructing 3D anatomical structures from freehand 2D transvaginal ultrasound sweeps, without requiring external tracking or learned pose estimators. Our method, TVGS, adapts the principles of Gaussian Splatting to the domain of ultrasound, introducing a slice-aware, differentiable rasterizer tailored to the unique physics and geometry of ultrasound imaging. We model anatomy as a collection of anisotropic 3D Gaussians and optimize their parameters directly from image-level supervision. To ensure robustness against irregular probe motion, we introduce a joint optimization scheme that refines slice poses alongside anatomical structure. The result is a compact, flexible, and memory-efficient volumetric representation that captures anatomical detail with high spatial fidelity. This work demonstrates that accurate 3D reconstruction from 2D ultrasound images can be achieved through purely computational means, offering a scalable alternative to conventional 3D systems and enabling new opportunities for AI-assisted analysis and diagnosis.
@@ 1 Introduction
Three-dimensional (3D) ultrasound (US) imaging plays an essential role in obstetrics and gynecology, providing clinicians with a spatial context that is crucial for diagnosis, surgical planning, and treatment monitoring [1, 11]. Transvaginal ultrasound (TVS) is widely used for high-resolution imaging of pelvic organs due to its affordability, safety, and minimally invasive nature, but extending 2D TVS to 3D imaging remains technically challenging. Existing 3D US systems typically rely on external tracking mechanisms or mechanically swept probes. However, while 3D ultrasound imaging is commonly performed both transabdominally and transvaginally, especially in obstetric applications such as fetal imaging, most commercial 3D systems are designed for abdominal or general freehand use, not specifically for transvaginal gynecological contexts [2, 22]. In addition to being physically cumbersome and difficult to integrate into routine clinical workflows, these systems often suffer from lower resolution compared to conventional 2D imaging. The reliance on mechanical scanning or external tracking can introduce motion artifacts, particularly in dynamic or less controlled scanning environments. Furthermore, the need for specialized hardware increases system complexity and cost, which hinder widespread adoption.
@@
In this work, we propose TVGS: an unsupervised method for 3D US reconstruction that is free from the need for any external tracking or prior learned pose estimation. Our method is specifically designed for the gynecological domain and operates on standard 2D TVS sweeps acquired freehand. It leverages recent advances in volumetric rendering, adapting Gaussian Splatting [12] to the constraints and physics of US imaging. Our approach uses sensorless geometric modeling of probe motion to establish spatial coherence between slices, and reconstructs anatomical structures using anisotropic Gaussian primitives optimized directly from US data.
Our framework represents anatomy as a collection of anisotropic Gaussians, optimized via a custom CUDA-optimized differentiable rasterizer. This continuous representation mitigates the aliasing and staircase artifacts inherent to discrete voxel-based methods, particularly in off-axis slices, while achieving convergence speeds significantly faster than implicit neural representations. Importantly, our method requires no calibration data and does not rely on training with large datasets.
@@
  • A CUDA-optimized differentiable rasterizer tailored for slice-based rendering under highly anisotropic sampling (dense in-plane, sparse through-plane).
  • A joint optimization strategy that refines slice poses alongside scene parameters, increasing robustness to irregular probe motion typical in freehand acquisition.
@@ 2 Related Work
Early systems using electromagnetic sensors [21] provided accurate positioning but are ill-suited for TVS due to bulk and patient discomfort [17]. Consequently, sensorless methods emerged, leveraging signal processing techniques like speckle decorrelation [3, 8, 13] and spatial correlation [9] to infer probe movement. While foundational, these approaches often struggle with noise and complex multi-axis motion. Later statistical and graph-based optimizations [20] improved robustness but remained too fragile for widespread clinical adoption.
@@
With the rise of deep learning, supervised learning methods for pose estimation gained popularity. These approaches train convolutional or recurrent neural networks to predict either inter-frame transformations or directly reconstruct 3D volumes. Early models demonstrated that CNNs could infer relative probe motion from US frames, enabling trackerless 3D reconstruction [18]. Later methods like $\text{DC}^{2}$-Net [10] introduced contrastive and temporal losses to improve trajectory consistency across sequences. Another noteworthy model is PLPPI, which incorporates domain-specific physics constraints into the learning pipeline, enabling improved handling of out-of-plane motion [4]. However, a shared limitation of these models is their dependency on ground truth data from tracked acquisitions for training and evaluation. In many clinical workflows—particularly for TVS—such tracking data is unavailable or infeasible to collect. Some recent works like RecON [15] explore online learning, updating models in real-time as new frames are acquired. While promising for reducing drift and latency, these methods still presuppose access to prior 3D volumes or strong priors, making them unsuitable for fully unsupervised clinical settings.
@@
In parallel, implicit representations like Neural Radiance Fields (NeRF) [16] and subsequent developments (e.g., ImplicitVol [22]) model 3D structures as continuous functions learned by neural networks. These models offer resolution independence and are memory-efficient but typically require extensive training and known camera poses—conditions that are generally not met in US unless a tracking device is present. Hybrid methods such as RapidVol [6] attempt to blend implicit and explicit approaches, using tri-planar maps refined by lightweight networks to accelerate training and inference. However, even these models typically assume access to 3D volumes for training or initialization.
@@
Beyond motion estimation, the choice of volumetric representation plays a central role in reconstruction quality and efficiency. Traditional explicit methods such as voxel grids offer high interpretability and easy integration into existing medical imaging pipelines. However, they suffer from cubic memory growth with resolution and are thus computationally intensive for fine anatomical detail. More recently, techniques like Gaussian Splatting have emerged as a powerful alternative. Originally developed for real-time rendering in computer vision [12], Gaussian Splatting represents scenes using parameterized Gaussian primitives whose locations, shapes, and radiance are optimized to match rendered views. UltraGauss [5] adapted this to fetal US imaging by incorporating anisotropic Gaussians and integrating pose predictions from QAERTS [19]. This approach improves rendering quality and speed but still depends on learned pose models, which hinders generalization and deployment in tracking-free environments.
@@
In contrast to these existing approaches, our method introduces the first application of Gaussian Splatting for 3D US that is entirely unsupervised and sensorless. We eliminate the need for pose estimation models, supervision, or tracking hardware by leveraging geometric assumptions on sweep motion and optimizing anisotropic Gaussian primitives directly from 2D TVS slices. This enables anatomically accurate, scalable 3D reconstruction using only standard 2D US equipment, significantly lowering the barrier to clinical deployment and advancing the capabilities of low-cost US imaging.
@@ 3.1 Synthetic Data
Figure 2: Synthetic ultrasound slices generated from a simulated uterus model using a renderer. Left: Sagittal view showing longitudinal anatomy. Right: Transversal view perpendicular to the sagittal plane. These slices can be sampled at arbitrary resolution.
@@ 3.2 Real Data
20 TVS sweeps were acquired from 10 patients (2 sweeps per patient) using a GEHC Voluson Expert 22 ultrasound machine. Table 1 summarizes the main parameters of the data. Each sweep represents either a sagittal or transversal view, with probe motion approximated as a smooth angular trajectory. An example image for both modalities can be seen in Fig. 3. Each patient provided written informed consent prior to participation, and all procedures adhered to strict ethical standards to ensure confidentiality and privacy. Since real data lacks tracked probe poses, we assume consistent frame intervals and smooth probe motion. We initialize the pose of each frame using equal angular spacing. This assumption mirrors the angular sampling strategy used in generating synthetic data. Other than resizing and padding images when needed, no additional transformations or data augmentation techniques are applied, ensuring that the intrinsic properties of the images are preserved. Our method assumes (i) temporally ordered slices with approximately uniform angular spacing, and (ii) minimal tissue deformation during the sweep. These assumptions are critical for effective sensorless reconstruction, and we analyze their impact in Section 4.
@@
Frames per second 10-20
Duration $\sim$10 seconds
@@
Original Resolution 1528 x 784
Age 41.9 (6.0)
Height [m] 1.67 (0.07)
Weight [kg] 68.49 (13.26)
@@
3.3 Test Set
For quantitative evaluation, we establish a distinct test set partition. Although the raw sweep lengths vary, we standardize our evaluation by subsampling 85 equidistant slices from each sweep. From these, we select 10 slices uniformly at random to serve as the test set and keep the remaining 75 for training (we analyze the impact of different slice counts in Section 5.3.3). Crucially, because the ultrasound slices represent thin, planar cross-sections with sparse angular spacing, these held-out test slices do not spatially overlap with any of the remaining slices used for training. Consequently, this evaluation setup is particularly challenging; it effectively constitutes a task of novel view synthesis (NVS), requiring the model to accurately infer anatomical structures in the gaps between training views without direct supervision. As we do not have accurate poses for NVS, we perform a short optimization of them before evaluation to factor out pose error, consistent with standard practices for pose estimation in neural rendering frameworks [14].
@@ 4.1 Overview: From Gaussian Splatting to Ultrasound Imaging
However, US imaging presents fundamental differences. Each image (or “slice”) corresponds to a narrow, planar cross-section of tissue, implying limited overlap between any two slices. In TVS, this limitation is even more pronounced as there is no overlap between slices in the sweep. As a result, there is no redundancy between views to help maintain scene consistency. Additionally, the physics of US introduces directionality and discontinuities that standard perspective projection does not capture. This stands in contrast to Gaussian Splatting for optical images, where we work with projections of the scene onto a camera sensor, inherently providing depth information. Since US does not have this projection component and instead captures direct cross-sections, approaches relying on projection geometry cannot be directly applied. Misaligned Gaussian primitives—especially in depth—can significantly affect the rendering, since each slice is extremely thin. To address these challenges, we adapt Gaussian Splatting, building a slice-aware differentiable rasterizer specifically designed for US data. Our framework accounts for the sparse and largely non-overlapping slices while retaining the strengths of the Gaussian formulation—compact, continuous scene representation and differentiability.
@@ 4.2 Slice-Based Differentiable Rasterization
Our rasterizer computes the intensity of each pixel in a slice by aggregating contributions from a set of 3D Gaussian primitives. We begin by describing the formulation in terms of its constituent steps. Each Gaussian $g$ is defined by a set of parameters:
@@
  • $\Sigma_{g}\in\mathbb{R}^{3\times 3}$: the covariance matrix describing its shape and orientation. It is factored in the same way as in the original GS formulation [12]: $\Sigma_{g}=(S_{g}R_{g})^{\top}(S_{g}R_{g})$, where $S_{g}$ is a diagonal scale matrix and $R_{g}$ is a rotation matrix derived from a unit quaternion,
@@
Each pixel $p$ in a US slice is associated with a known 3D coordinate $\mathbf{c}_{p}\in\mathbb{R}^{3}$, obtained via slice pose modeling as described in Section 3.2 and the location of the pixel in the slice. Notably, $\mathbf{c}_{p}$ is a differentiable function of the slice’s rigid body transformation parameters. Consequently, gradients computed at the pixel level during the backward pass flow not only to the Gaussian parameters but also to the extrinsic matrix of the slice, enabling the correction of misalignment errors. For each Gaussian $g$, we compute the offset vector from the Gaussian center to the pixel coordinate:
@@
This formulation defines each pixel as the sum of soft and spatially-varying influences from 3D Gaussians. By using Gaussian functions instead of hard surface projections, the method can represent continuous, soft-tissue-like structures and remains differentiable for optimization. It is especially well-suited for sparse data with anisotropic spatial sampling such as US slices, where direct projection is unreliable and information across slices is limited.
To enable efficient end-to-end optimization of this representation, we implement the rasterizer as a high-performance CUDA kernel. The custom nature of this slice-accumulation process precludes the use of standard automatic differentiation frameworks like PyTorch’s autograd. Instead, we analytically derive the gradients for all Gaussian parameters and implement them directly within the kernel using the chain rule. This manual backward pass also avoids heavy automatic differentiation graphs, resulting in significant memory savings.